OpenStudy (anonymous):

13x-2y=70 2x-2y=4 Can someone please help me?

OpenStudy (anonymous):

so you are trying to find y and x correct?

OpenStudy (loser66):

your trying?

OpenStudy (anonymous):

wait are you talking to me or @WonderlandKy

OpenStudy (anonymous):

i have the correct answer @WonderlandKy but i'm not going to give it to you you have to try to solve it too

OpenStudy (loser66):

@shinebrightlikeadimon of course, the problem who need trying is the Asker, not you

OpenStudy (loser66):

*person, not problem

OpenStudy (anonymous):

Yes, I'm trying to find x and y. I know that you plug numbers in but I've done big numbers and small and can't seem to find numbers that fit both.

OpenStudy (anonymous):

so your plunging in numbers and guessing?

OpenStudy (anonymous):

and yah i feel the same way like i want people to teach me how to do it rather then give me the answer @Loser66

OpenStudy (anonymous):

13x-2y=70 2x-2y=4 1) 13x-2y=70 --> rearrange to get 2y = 13x-70 Substitute that value for 2y into the 2nd equation: 2x-2y=4 2x - (13x-70) = 4 2x - 13x + 70 = 4 -11x = -66 x = 6 2) Now substitute this value for x into EITHER of the two original equations to solve for y... 13x-2y = 70 13(6) - 2y = 70 78 - 2y = 70 -2y = -8 y = 4 Solution set: \[\left\{ 6,4 \right\}\]

OpenStudy (anonymous):

i agree with your answer ^^ but you should have at least let the asker try

OpenStudy (anonymous):

I appreciate the input @shinebrightlikeadimon. However, having seen an example worked out fully, he/she can now use what they've learned and apply this method to additional problems going forward. Good luck, @WonderlandKy!

OpenStudy (anonymous):

sorry if i offended you but the person can just look at what you wrote last and not care about the rest

OpenStudy (anonymous):

Agreed, and they would be harming noone but themselves by doing that.

OpenStudy (anonymous):

...and no offense taken. :-)

OpenStudy (anonymous):

that's true and ok